# Title keyword 5 for citation Title for citation Title for citation

### v2

@@ 4.2.1 Distributed Intelligence and Federated Learning
Multi-Agent Reinforcement Learning: This branch of reinforcement learning studies the behavior of multiple learning agents coexisting in a shared environment. Each agent pursues its reward and acts to advance its interests. The primary goal is to achieve cooperative and win-win strategy learning among multiple distributed nodes, suitable for multi-agent collaboration or competition scenarios. 6G networks must support diverse AI application scenarios, such as smart cities and intelligent transportation. Multi-agent reinforcement learning (MARL), with its flexible learning and decision-making capabilities, can offer personalized and optimized solutions across various applications. By learning the characteristics of their environment and tasks, agents can autonomously optimize and adjust to meet different application needs, enhancing user experience.
MARL has been widely applied in practice, making progress in smart cities, intelligent transportation, smart manufacturing, personalized services, and unmanned systems. The study in [190] presents, for the first time, a fully scalable and decentralized MARL algorithm for adaptive traffic signal control in complex urban traffic networks. The authors of [191] propose a novel multi-agent recurrent deep deterministic policy gradient algorithm based on a deep deterministic policy gradient (DDPG) algorithm for traffic light control in vehicular networks. The authors of [192] present new cyber-physical integration in smart factories for online scheduling of low-volume-high-mix orders. The authors of [193] propose a novel multi-agent deep reinforcement learning scheme to solve the multichannel access and task offloading problem in MEC-enabled industry 4.0. The authors of [194] first apply deep reinforcement learning to the problem of forming long-term driving strategies. The authors of [195] propose using Partially Observable Markov Games to formulate the connected autonomous driving problems with realistic assumptions.
Split Learning: This technique separates machine learning model architecture between clients and servers, allowing joint model training without disclosing raw data. High-computation tasks are assigned to powerful central nodes, while shallower data processing layers are kept at data-holding terminals. Split learning offers better privacy than FL and reduces the processing load on resource-limited devices. Its applications focus on healthcare, where lightweight devices train intelligent models based on users’ private health data [196].
The following are the future research directions of distributed intelligence [197]:
(1) Future distributed networks will focus more on improving performance and reliability. More effective error correction and retransmission mechanisms can be adopted to address data loss and errors in unreliable communication environments, such as advanced error-correcting codes and adaptive retransmission strategies, to reduce error rates and enhance communication reliability.
(2) As data breaches and network attacks become more prominent, future distributed networks will prioritize security and privacy protection. Enhanced data encryption, identity authentication, and access control measures will protect user data from unauthorized access and tampering.
@@ 4.2.2 Digital Twins for 6G Network native AI
In terms of management and orchestration services, DT is a key enabling technology that enables more efficient and accurate allocation of resources and system optimization for AI services. In B5G/6G communication systems, the rapid growth in user demand and the continuous evolution of technical standards have posed significant challenges to resource allocation and system optimization. Traditional communication systems, which rely on historical data and static models, struggle to adapt swiftly to dynamic network changes, making real-time adjustments difficult. These limitations hinder the adaptability, reliability, and flexibility of communication networks. DT technology has been introduced into the communication domain to address these challenges, offering a flexible, efficient, and precise approach to resource optimization. Adopting DT not only enhances system agility and performance but also introduces innovative models for network operation, significantly reducing research and development costs and risks [198]. Figure 19 depicts the relationship between 6G physical infrastructure and DT for AI.
@@
The DT platform employs physical models and data-driven machine learning algorithms to construct high-fidelity virtual replicas that comprehensively simulate physical entities’ operational state and dynamic changes. These virtual replicas accurately reflect the current state and leverage simulation techniques to predict future system behaviors or potential anomalies. With bidirectional data interaction, physical systems and virtual replicas achieve real-time data synchronization. The DT analyzes large-scale data intelligently, generates optimized decisions, and swiftly relays these to the physical system. For instance, when a network traffic surge is anticipated, the DT can proactively adjust load distribution strategies, optimize spectrum and power resources, and alleviate stress on network devices.
@@
The 6G network aims to achieve ultra-high capacity, ultra-short-range communication, beyond-best-effort and high-precision communication, and integrate multiple communication types [202]. However, this ambitious vision introduces security, spectrum efficiency, intelligence, energy efficiency, and affordability challenges. The emergence of DT technology offers promising opportunities to address these challenges. DT provides a virtualized counterpart for 6G networks, enabling comprehensive network traffic monitoring and analysis. By collecting traffic data across the entire network, DT utilizes data analytics to proactively uncover traffic patterns and detect anomalies. Leveraging feedback from the virtualized network, 6G systems can enhance their security by preparing for potential threats in advance. Additionally, DT enables the automation of demand identification and service provisioning by analyzing communication data to discern usage patterns and rules. Predictive insights into communication demands allow 6G networks to reserve resources, such as spectrum, in anticipation of future needs. Moreover, the integration of DT technology empowers 6G to support innovative services, including augmented reality (AR), virtual reality (VR), and autonomous driving. By addressing key challenges in security, spectrum efficiency, intelligence, energy efficiency, and customization, DT technology redefines and accelerates the development of 6G networks [203].
@@
Significant research has integrated DT technology with 6G networks to enhance network performance. For instance, DT has been proposed for managing metasurface reflectors in 6G terahertz communications [204]. By modeling, predicting, and controlling the propagation characteristics of indoor signals, DT maximizes the system’s terahertz signal-to-noise ratio (SNR). Additionally, Lu et al. introduced the concept of DT Wireless Networks (DTWN) by integrating DT into wireless networks. This approach shifts real-time data processing and computation to the edge plane and leverages DT to mitigate the unreliability of long-distance communications between end users and edge servers in 6G networks [205]. Overall, the deep integration of DT with 6G effectively bridges physical systems and digital spaces, providing robust technical support for efficient and reliable wireless communication.
@@
One of the critical capabilities of 6G networks will be supporting massive-scale terminal devices that generate extensive data traffic. In this context, DT can optimize edge resources to meet user demands better. For example, the authors of [206] designed a DT model for edge computing in 6G wireless networks. By incorporating mobile users’ dynamic network state and DT, the study addressed edge association problems using deep reinforcement learning and transfer learning, successfully reducing average system latency and improving resource utilization. Similarly, the authors of [207] proposed a DT-assisted task offloading method, creating DTs for all device states to achieve lower latency and power consumption in the network. Furthermore, the authors of [208] investigated dynamic resource allocation mechanisms for DT-based services in the 6G Internet of Things (IoT). The study improved resource allocation efficiency by establishing service function chains with DTs and employing a collective reinforcement learning method.
@@
In addition, Digital Twin Channel (DTC) is also a research direction of DT for 6G Network native AI. DTC is proposed as a digital virtual mapping of a wireless channel that reflects the entire process of channel fading states and variations in the physical world [209]. DTC is envisaged to continually interact with the wireless communications system (WCS) channel requirements to support the system’s ability to adapt to complex communication environments actively. It can be applied in various typical 6G usage scenarios, offering significant performance enhancements in data rate, latency, spectral efficiency, energy efficiency, and link reliability [210]. To boost the capabilities of the DTC for the predictive 6G network, a novel DTC implementation platform is proposed, including data acquisition, environment sensing and reconstruction module to build the relationship between environment and channel for channel prediction, communication decision and interaction [211, 212]. A cluster-nuclei (CN) based channel model is proposed to analyze the mapping relationship between clusters and scatterers in the propagation environment, enabling channel reconstruction across various scenarios [213]. In addition, a radio environment knowledge pool (REKP) is proposed to serve as a specific enabler for DTC. Based on the principles of REKP, the electromagnetic wave property inspired radio environment knowledge (REK) construction method is proposed, which fills the gap in the interpretable mathematical representation of the relationship between adequate environment information and channel [214].
@@
The ITU documentation defines the requirements and architecture of DT Networks (DTNs), emphasizing their core role as virtual representations of physical networks. By leveraging data collection, real-time interaction mapping, unified data models, and standardized interfaces, DTNs enable physical network analysis, diagnosis, simulation, and control. The ITU proposal introduces a ”three-layer, three-domain, double closed-loop” architecture design, highlighting applications in complex network operations, optimization, innovation, and security strategy drills. DTNs’ functional requirements include data management, modeling, and lifecycle management, while service requirements focus on compatibility, scalability, and privacy, further supporting the intelligent, automated, and secure development of future networks.
@@
DT technology is pivotal in developing 6G systems and various communication domains. By enabling precise modeling and real-time interaction with physical entities, DT provides practical solutions for optimizing resource allocation, enhancing communication security, and improving service management. For instance, DT optimizes computation offloading in industrial IoT in ultra-reliable, low-latency communication links. Integrating DT with blockchain and FL technologies in wireless communication strengthens system security and privacy protection. In the vehicular network domain, DT leverages traffic prediction, edge computing, and load-balancing techniques to enhance real-time data processing and overall system efficiency. DT is driving innovation and development across multiple dimensions of communication systems, demonstrating its transformative potential. The application of DT technology in communication systems is currently experiencing rapid development, showcasing significant potential across various domains, particularly in the research and development of 5G and the emerging 6G networks. In industrial IoT environments, DT technology plays a crucial role. For instance, the authors of [215] address minimizing computation offloading delays in ultra-reliable low-latency communication (URLLC) links by leveraging DT-enabled wireless edge networks. The study focuses on modeling the computational capabilities of edge servers and optimizing system-wide resource allocation.
@@
To enable wireless communication systems based on proactive online learning, the authors of [216] introduce a DT design framework. This framework considers aspects such as twin object access, security and privacy, and air interface design, while also discussing the challenges and opportunities of DT in wireless systems. Additionally, the authors of [217] propose a blockchain-empowered FL framework to enhance communication security and data privacy within DT-enabled edge networks. The study incorporates asynchronous aggregation mechanisms and DT-driven reinforcement learning to optimize user scheduling and spectrum resource allocation.
@@
DT technology has also seen applications in vehicular networks. The authors of [218] present a DT-assisted real-time traffic data prediction method capable of delivering accurate short-term forecasts for traffic flow and speed. The authors of [219] propose an innovative vehicular edge computing network combining DT and multi-agent learning. This approach helps uncover potential matches for edge services among large-scale vehicles, reducing the complexity of service management. Similarly, the authors of [220] introduce a DT-based load-balancing prediction model for autonomous vehicles, which accurately predicts road network conditions while ensuring high data transmission security. Furthermore, the authors of [221] explore the use of DT in vehicular edge computing, developing models that reflect the real-time state of the vehicular environment. The study introduces two metrics, QDT and CDT, to evaluate the quality and cost of DT models for edge nodes, optimizing system performance and reducing costs. These studies highlight DT’s transformative potential across communication systems, from resource optimization in industrial IoT to service management and predictive modeling in vehicular networks.
@@
The application of DT technology in communication systems has made initial progress but is still in the exploratory stage. In the future, DT technology needs to improve model construction methods and accuracy further, optimize real-time performance and latency, enhance security and privacy, and reduce resource consumption and costs. With AI and machine learning technology advancements, DT is expected to provide more intelligent network operation and maintenance solutions, achieving more efficient fault prediction, network optimization, and service quality management.

### v3

@@ 3.2.1 Distributed Intelligence and Federated Learning
Multi-Agent Reinforcement Learning: This branch of RL studies the behavior of multiple learning agents coexisting in a shared environment. Each agent pursues its reward and acts to advance its interests. The primary goal is to achieve cooperative and win-win strategy learning among multiple distributed nodes, suitable for multi-agent collaboration or competition scenarios. 6G networks must support diverse AI application scenarios, such as smart cities and intelligent transportation. The multi-agent reinforcement learning (MARL), with its flexible learning and decision-making capabilities, can offer personalized and optimized solutions across various applications. By learning the characteristics of their environment and tasks, agents can autonomously optimize and adjust to meet different application needs, enhancing user experience.
The authors of [154] proposed an intelligent and decentralized MARL algorithm based on advantage actor critic to solve large-scale traffic signal control problems, providing an effective solution to the non-stationary problem of congestion relief in intelligent transportation systems. Similarly, for the problem of controlling traffic lights in cities, the authors of [155] proposed a new multi-agent recurrent DDPG algorithm, which adjusts the phase of traffic signals in real time through Internet of Vehicles data, effectively reducing traffic congestion under complex road conditions and partially observable environments. In addition, in the intelligent transportation, the authors of [156] proposed the use of partially observable markov game to formulate connected autonomous driving problems with realistic assumptions, and developed a multi-agent interconnected autonomous driving learning platform, providing a new practical basis for the development of autonomous driving technology. In [157], the authors focused on achieving collaborative computing between multiple layers of low earth orbit satellites. By establishing a sequential model for task processing and obtaining the connection relationship between satellites in different orbits, a real-time offloading decision model was constructed and the use of MARL combined with penalty function method was proposed to optimize the average performance of the system under continuous time scales, providing a new solution for improving the system performance of satellites supporting edge computing under non-uniform loads. In [158], the authors introduced a MARL-based online production scheduling method in smart factories. Through distributed computing architecture and compound reward functions, they effectively proved that the AI-based online solution is superior to the conventional solution in production scheduling performance, and shows stronger adaptability and robustness under uncertain situations such as order changes and machine failures. The authors of [159] proposed a novel multi-agent DRL scheme to solve the multi-channel access and task offloading problems in MEC-enabled Industry 4.0 and improved the computational efficiency and resource utilization of industrial systems.
MARL is suitable for signal control and autonomous driving applications in intelligent transportation, and is expected to become a promising solution in collaborative computing between satellites and production scheduling in intelligent industry, resource utilization optimization of task offloading, etc. However, it still has shortcomings such as high algorithm complexity and lack of data privacy protection.
Split Learning: Considering that devices with limited resources cannot support complex DL models and thus cannot support collaboration in the form of FL [160], the industry has derived split learning (SL) technology. SL splits the ML model and allows different parts of the model to be distributed between the client and the server, while allowing collaboration for joint model training without leaking the original data [161]. High computing tasks are assigned to powerful central nodes, while shallower data processing layers are retained at the terminal where the data is stored. This can make ML easier to use for resource-constrained devices. As an alternative/complementary method to FL, SL can provide better privacy protection and reduce the processing load of resource-constrained devices, and has the potential to become one of the key AI technologies for 6G distributed intelligence.
In the field of healthcare, the development of DL is constrained by the lack of data samples and data privacy protection. Based on this, the authors of [162, 163] introduced SL into healthcare applications to solve the problem of different entities collaboratively training DL models without sharing sensitive raw data, and achieved good experimental results on the model performance based on medical datasets. Similarly, in [164], the authors studied the use of SL for multi-institutional collaborative training of DL models, and verified that SL can achieve better client efficiency and more reliable privacy protection. Considering that the sequential training method of original SL may cause model overfitting or slow convergence, the authors of [165] proposed a parallel SL method, which solved the overfitting problem in sequential SL through training batch adaptive adjustment and node layer synchronization mechanism, and improved the training effect while ensuring privacy. In [166], the authors introduced SL into wireless networks to realize the utilization of a large amount of terminal privacy data in IoT. Considering that the sequential training method leads to large delays, they proposed a cluster-based parallel training method, and simultaneously optimized the joint cutting layer selection, device clustering and spectrum allocation to effectively reduce the training delay.
SL is suitable for multi-institutional collaborative training of DL models in healthcare with strict privacy requirements, and is also suitable for training efficient DL models in IoT. However, it has the disadvantages of being easily affected by data heterogeneity, complex training methods, overfitting of models, slow convergence, or even failure to converge.
The above are some typical representative technologies in distributed intelligence. Different technologies have different application scenarios and scopes of application. For FL, it is more suitable for efficiently implementing AI services in scenarios with data privacy requirements. And FL is widely applicable to various emerging intelligent demand applications of 6G owing to its compatibility with other intelligent learning paradigms. For MARL, it is more in line with the needs of online learning in complex environments and completes AI tasks in an observable and interactive environment. For separation learning, it is suitable for realizing collaborative learning of DL models in resource-constrained devices, and can adapt to different levels of resource constraints through multi-level model splitting.
With the development of the networks and the advancement of ML technology, future distributed intelligence may also evolve towards other research directions. (1) To develop more reliable distributed intelligence, more effective error correction and retransmission mechanisms can be adopted to address data loss and errors in unreliable communication environments, such as advanced error correction codes and adaptive retransmission strategies to reduce the error rate and improve communication reliability. (2) To develop more secure distributed intelligence, more secure protocols and privacy protection mechanisms should be further considered. Enhanced data encryption, identity authentication and access control measures can be employed to protect user data from unauthorized access and tampering. In addition, there are some frontier directions that can be explored[125, 167], such as improving the communication efficiency of data/model parameter exchange, enhancing the energy efficiency of learning algorithms, and improving fairness through the incentive mechanisms for participating devices.
@@ 3.2.2 Digital Twins for 6G Network native AI
In terms of management and orchestration services, DT is a key enabling technology that enables more efficient and accurate allocation of resources and system optimization for AI services. In B5G/6G communication systems, the rapid growth in user demand and the continuous evolution of technical standards have posed significant challenges to resource allocation and system optimization. conventional communication systems, which rely on historical data and static models, struggle to adapt swiftly to dynamic network changes, making real-time adjustments difficult. These limitations hinder the adaptability, reliability, and flexibility of communication networks. DT technology has been introduced into the communication domain to address these challenges, offering a flexible, efficient, and precise approach to resource optimization. Adopting DT not only enhances system agility and performance but also introduces innovative models for network operation, significantly reducing research and development costs and risks [168]. Figure 17 depicts the relationship between 6G physical infrastructure and DT for AI.
@@
The DT platform employs physical models and data-driven ML algorithms to construct high-fidelity virtual replicas that comprehensively simulate physical entities’ operational state and dynamic changes. These virtual replicas accurately reflect the current state and leverage simulation techniques to predict future system behaviors or potential anomalies. With bidirectional data interaction, physical systems and virtual replicas achieve real-time data synchronization. The DT analyzes large-scale data intelligently, generates optimized decisions, and swiftly relays these to the physical system. For instance, when a network traffic surge is anticipated, the DT can proactively adjust load distribution strategies, optimize spectrum and power resources, and alleviate stress on network devices.
@@
The 6G network aims to achieve ultra-high capacity, ultra-short-range communication, beyond-best-effort and high-precision communication, and integrate multiple communication types [172]. However, this ambitious vision introduces security, spectrum efficiency, intelligence, energy efficiency, and affordability challenges. The emergence of DT technology offers promising opportunities to address these challenges. DT provides a virtualized counterpart for 6G networks, enabling comprehensive network traffic monitoring and analysis. By collecting traffic data across the entire network, DT utilizes data analytics to proactively uncover traffic patterns and detect anomalies. Leveraging feedback from the virtualized network, 6G systems can enhance their security by preparing for potential threats in advance. Additionally, DT enables the automation of demand identification and service provisioning by analyzing communication data to discern usage patterns and rules. Predictive insights into communication demands allow 6G networks to reserve resources, such as spectrum, in anticipation of future needs. Moreover, the integration of DT technology empowers 6G to support innovative services, including AR/VR, and autonomous driving. By addressing key challenges in security, spectrum efficiency, intelligence, energy efficiency, and customization, DT technology redefines and accelerates the development of 6G networks [173].
@@
Significant research has integrated DT technology with 6G networks to enhance network performance. For instance, DT has been proposed for managing metasurface reflectors in 6G terahertz communications [174]. By modeling, predicting, and controlling the propagation characteristics of indoor signals, DT maximizes the system’s terahertz SNR. Additionally, Lu et al. introduced the concept of DT wireless networks by integrating DT into wireless networks. This approach shifts real-time data processing and computation to the edge plane and leverages DT to mitigate the unreliability of long-distance communications between end users and edge servers in 6G networks [175]. Overall, the deep integration of DT with 6G effectively bridges physical systems and digital spaces, providing robust technical support for efficient and reliable wireless communication.
@@
One of the critical capabilities of 6G networks will be supporting massive-scale terminal devices that generate extensive data traffic. In this context, DT can optimize edge resources to meet user demands better. For example, the authors of [176] designed a DT model for edge computing in 6G wireless networks. By incorporating mobile users’ dynamic network state and DT, the study addressed edge association problems using DRL and transfer learning, successfully reducing average system latency and improving resource utilization. Similarly, the authors of [177] proposed a DT-assisted task offloading method, creating DTs for all device states to achieve lower latency and power consumption in the network. Furthermore, the authors of [178] investigated dynamic resource allocation mechanisms for DT-based services in the 6G IoT. The study improved resource allocation efficiency by establishing service function chains with DTs and employing a collective RL method.
@@
In addition, DT channel is also a research direction of DT for 6G Network native AI. DTC is proposed as a digital virtual mapping of a wireless channel that reflects the entire process of channel fading states and variations in the physical world [179]. DTC is envisaged to continually interact with the wireless communications system channel requirements to support the system’s ability to adapt to complex communication environments actively. It can be applied in various typical 6G usage scenarios, offering significant performance enhancements in data rate, latency, spectral efficiency, energy efficiency, and link reliability [180]. To boost the capabilities of the DTC for the predictive 6G network, a novel DTC implementation platform is proposed, including data acquisition, environment sensing and reconstruction module to build the relationship between environment and channel for channel prediction, communication decision and interaction [181, 182]. A cluster-nuclei based channel model is proposed to analyze the mapping relationship between clusters and scatterers in the propagation environment, enabling channel reconstruction across various scenarios [183]. In addition, a radio environment knowledge (REK) pool is proposed to serve as a specific enabler for DTC. Based on the principles of REK pool, the electromagnetic wave property inspired REK construction method is proposed, which fills the gap in the interpretable mathematical representation of the relationship between adequate environment information and channel [184].
@@
The ITU documentation defines the requirements and architecture of DT networks, emphasizing their core role as virtual representations of physical networks. By leveraging data collection, real-time interaction mapping, unified data models, and standardized interfaces, DTNs enable physical network analysis, diagnosis, simulation, and control. The ITU proposal introduces a “three-layer, three-domain, double closed-loop” architecture design, highlighting applications in complex network operations, optimization, innovation, and security strategy drills. DTNs’ functional requirements include data management, modeling, and lifecycle management, while service requirements focus on compatibility, scalability, and privacy, further supporting the intelligent, automated, and secure development of future networks.
@@
DT technology is pivotal in developing 6G systems and various communication domains. By enabling precise modeling and real-time interaction with physical entities, DT provides practical solutions for optimizing resource allocation, enhancing communication security, and improving service management. For instance, DT optimizes computation offloading in industrial IoT in uRLLC links. Integrating DT with blockchain and FL technologies in wireless communication strengthens system security and privacy protection. In the vehicular network domain, DT leverages traffic prediction, edge computing, and load-balancing techniques to enhance real-time data processing and overall system efficiency. DT is driving innovation and development across multiple dimensions of communication systems, demonstrating its transformative potential. The application of DT technology in communication systems is currently experiencing rapid development, showcasing significant potential across various domains, particularly in the research and development of 5G and the emerging 6G networks. In industrial IoT environments, DT technology plays a crucial role. For instance, the authors of [185] address minimizing computation offloading delays in uRLLC links by leveraging DT-enabled wireless edge networks. The study focuses on modeling the computational capabilities of edge servers and optimizing system-wide resource allocation.
@@
To enable wireless communication systems based on proactive online learning, the authors of [186] introduce a DT design framework. This framework considers aspects such as twin object access, security and privacy, and air interface design, while also discussing the challenges and opportunities of DT in wireless systems. Additionally, the authors of [187] propose a blockchain-empowered FL framework to enhance communication security and data privacy within DT-enabled edge networks. The study incorporates asynchronous aggregation mechanisms and DT-driven RL to optimize user scheduling and spectrum resource allocation.
@@
DT technology has also seen applications in vehicular networks. The authors of [188] present a DT-assisted real-time traffic data prediction method capable of delivering accurate short-term forecasts for traffic flow and speed. The authors of [189] propose an innovative vehicular edge computing network combining DT and multi-agent learning. This approach helps uncover potential matches for edge services among large-scale vehicles, reducing the complexity of service management. Similarly, the authors of [190] introduce a DT-based load-balancing prediction model for autonomous vehicles, which accurately predicts road network conditions while ensuring high data transmission security. Furthermore, the authors of [191] explore the use of DT in vehicular edge computing, developing models that reflect the real-time state of the vehicular environment. The study introduces two metrics, Quantum DT and Cognitive DT, to evaluate the quality and cost of DT models for edge nodes, optimizing system performance and reducing costs. These studies highlight DT’s transformative potential across communication systems, from resource optimization in industrial IoT to service management and predictive modeling in vehicular networks.
@@
The application of DT technology in communication systems has made initial progress but is still in the exploratory stage. In the future, DT technology needs to improve model construction methods and accuracy further, optimize real-time performance and latency, enhance security and privacy, and reduce resource consumption and costs. With AI and ML technology advancements, DT is expected to provide more intelligent network O&M solutions, achieving more efficient fault prediction, network optimization, and service quality management.